OpenStudy (anonymous):

Please help,will fan and medal !! write the units that should appear in the answer to each of the following: a.g/ml*ml= B.g/ml*g*ml/cm^3 c.64.8642g+8.4g D.0.0042 ml*864g*95 ml/g=

OpenStudy (cuanchi):

a. g b.? please use parenthesis to indicate if the units are in the numerator or denominator. (g/ml)* g*(ml/cm^3) = g^2/cm^3 (doesnt make too much sense) (g/ml*g)(ml/cm^3) = 1/cm^3 = cm^-3 c. g d.ml^2 (doesn't make too much sense)

OpenStudy (anonymous):

@Cuanchi Yeah thats all it gives me chemeistry is crazyy but ehhh..:))